The velocity of a physical system is given by v=[{p+1/n}/x]½.

Where p is the pressure, find the dimensions of n and x.


1
Expert's answer
2021-03-24T20:06:13-0400

Solution.

"v=\sqrt{\frac{p+\frac{1}{n}}{x}};" "[v]=1ms^{-1};" "p=\frac{F}{s};" "[p]=1Nm^{-2};"

"[F]=1N=1kgms^{-2};"

"[\frac{1}{n}]=1Nm^{-2};\implies [n]=1N^{-1}m^{2}=1kg^{-1}m^{-1}s^2m^{2}=
1kg^{-1}ms^2;" "x=\frac{p+\frac{1}{n}}{v^2};" "[x]=\frac{1Nm^{-2}}{1m^2s^{-2}}=1Nm^{-4}s^2=1kgms^{-2}m^{-4}s^{2}=1kgm^{-3};"

Answer: "[n]=1kg^{-1}ms^2; [x]=1kgm^{-3}."